# Mapping fNIRS Signals to Agent Performance: Toward Reinforcement Learning from Neural Feedback

### v1

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) is a methodology that aligns agent behavior with human preferences by integrating human feedback into the agent’s training process. We introduce a possible framework that employs passive Brain-Computer Interfaces (BCI) to guide agent training from implicit neural signals. We present and release a novel dataset of functional near-infrared spectroscopy (fNIRS) recordings collected from 25 human participants across three domains: a Pick-and-Place Robot, Lunar Lander, and Flappy Bird. We train classifiers to predict levels of agent performance (optimal, sub-optimal, or worst-case) from windows of preprocessed fNIRS feature vectors, achieving an average F1 score of 67% for binary classification and 46% for multi-class models averaged across conditions and domains. We also train regressors to predict the degree of deviation between an agent’s chosen action and a set of near-optimal policies, providing a continuous measure of performance. We evaluate cross-subject generalization and demonstrate that fine-tuning pre-trained models with a small sample of subject-specific data increases average F1 scores by 17% and 41% for binary and multi-class models, respectively. Our work demonstrates that mapping implicit fNIRS signals to agent performance is feasible and can be improved, laying the foundation for future brain-driven RLHF systems.
@@
Codehttps://github.com/your-profile/NeuroLoop-Classification
@@ Introduction
Recent advances in Reinforcement Learning from Human Feedback (RLHF) have become crucial for training and fine-tuning state-of-the-art systems  (wu2023finegrained; mosqueiraSurvey). Specifically, RLHF has seen significant success by addressing certain limitations that plague general Reinforcement Learning (RL) techniques. Integrating human feedback has been incredibly beneficial for aligning agent behavior with human preferences. However, many feedback methods require active participation and/or expert demonstration in order for evaluative feedback to be most useful (christiano; liRLsurveyRLHF; retzlaff2024human). In addition, most approaches are limited to explicit feedback, which may lack insight into the nuances of human decision-making and internal assessments. Techniques that retrieve richer feedback may require tasks that impose greater cognitive effort on the user (lindner2022humans). These drawbacks can be inconvenient for the evaluator or limit accessibility to certain user groups. They may also contribute to higher mental workload and unnatural, shallow feedback (casper2023open). These limitations may make it difficult to develop adaptive AI technologies that fully align with human preferences while being available to all user groups.
Figure 1: NEURO-LOOP: A high-level diagram of the proposed NEURO-LOOP framework. This paper addresses the neural classification problem with the intent to apply our trained models to the proposed pipeline.
Implicit human feedback systems present a valuable alternative by allowing agents to learn from natural human responses. Modern methods focus on facial expression or gesture classification to adapt agent behavior at little to no additional cognitive cost to the teacher (cui2021empathic). However, some of these procedures still require explicit instruction and conscious physical effort to adjust gestures and expressions. Further, reliance on private training data poses challenges for data collection and distribution.
To address these limitations, we propose NEURO-LOOP, a fully implicit neural feedback framework that leverages fNIRS signals to directly guide agents’ training using RLHF techniques. This framework employs passive Brain-Computer Interfaces (BCI) to align agent behavior with the user intent through passive observation. We employ a functional near-infrared spectroscopy (fNIRS) device to record brain activity. fNIRS is a non-invasive neuro-imaging tool that can track sustained cognitive states such as mental workload, attention, and decision-making. This technology measures the hemodynamic response, or change in blood flow, in the prefrontal cortex (PfC) over time.
This paper addresses the neural classification problem as a critical first step toward integrating implicit neural signals into RLHF frameworks. Our key contributions are outlined below:
@@
  • We propose that machine learning can distinguish multiple levels of agent performance, extending past binary classification and enabling finer-grained feedback while maintaining low cognitive workload for users.
  • We evaluate model performance and generalization between participants, suggesting that cross-participant transfer remains a challenge, but subject-specific calibration using limited data is feasible and improves performance.
We further compare model performance and user workload across both passive (observation-only) and active (physical demonstration) conditions, establishing a benchmark for future research in brain-informed RLHF.
@@
Reinforcement Learning from Human Feedback allows agents to learn from human preferences through a range of approaches. These methods can be described as having either implicit or explicit feedback mechanisms.
@@
Explicit Feedback: Explicit feedback is the most common modality used to shape an agent’s policy to align with human expectations. Methods in this category often provide binary or scalar evaluative signals, such as preference labels or human-provided rewards, to shape the agent’s policy or learning process (griffith; KnoxStone2011AugmentingRL). Past research notes that evaluative metrics with wider ranges have greater success during training (aabl). These interactions are often shorter-term due to increased cognitive workload and sustained physical demand. Imitation learning is a related approach that often requires expert demonstration or active participation for the agent to learn from user examples effectively (imitation).
@@
Implicit Feedback: Implicit feedback leverages unobtrusive human responses to enhance reinforcement learning agent behavior. These feedback techniques often include gaze tracking (Veeriah2016FaceValuing), expression recognition (Arakawa2018DQNTAMERHR), or gesture classification (cui2021empathic) to shape agent policies. One drawback is that some of these procedures require explicit instruction, as participants must consciously and physically adjust their gestures or expressions to provide meaningful feedback. We propose using neural signals to communicate evaluative feedback directly with minimal instruction.
@@
Figure 2: Setup: Participants sat 24 inches away from a computer screen. The fNIRS device is a headband that shines pulsating infrared light into the PfC to detect changes in blood flow.
@@
Figure 3: Training Paradigms: Diagram of the three training paradigms and Experimental Protocol. a) Single-subject models are trained on a set of data from one participant and are evaluated using withheld data from the same participant. b) Multi-subject models are trained on a set of participants and are evaluated using withheld data from the same set. c) Fine-tuned models are multi-subject models calibrated with a fraction of a target participant’s data.
@@
ErrPs for Human-Robot Interaction: Error-related potentials (ErrPs) are neural signals that occur when the brain perceives an error. These signals, recorded via EEG, have been explored as a method for improving human-robot interaction. Previous research has applied ErrP signals to intervene in robotic decision-making by providing a binary indicator of error or dissatisfaction. Some work has even integrated ErrPs into RLHF algorithms, demonstrating improved training efficiency  (vukelic_combining_2023; agarwal_humanloop_2020; xu_accelerating_2021). EEG signals are often transient and susceptible to artifacts (rihet_robot_2024). We explore the use of fNIRS as a complementary signal for brain-driven RLHF systems. Compared to other neuro-imaging devices, fNIRS produces signals that offer more tolerance to physical movement, greater portability, and higher spatial resolution (Pinti2018Noise). Using the fNIRS system described by Giles (giles), we collected phase and intensity signals, making fNIRS suitable for long-horizon naturalistic settings outside the laboratory.
@@
fNIRS for Human-Centered Technologies: Recent advancements are making it possible for brain signals to be used in human-centered technologies (humancenteredfnirs). When applied to adaptive frameworks, fNIRS data often monitors cognitive workload, allowing systems to adjust to user strain. Past work has explored applications like unmanned aerial vehicles (UAV) (afergan_dynamic_2014), and human-robot systems (solovey_brainput_2012; roy_how_2020).
@@
Motivation: fNIRS is quite under-explored in reinforcement learning and human-in-the-loop machine learning. fNIRS also offers beneficial characteristics and features that EEG generally lacks. Prior research has correlated fNIRS signals to reward-based gameplay events like video game performance (VideoGamefNIRS2018) and physical game performance (fNIRSRewardDanceSimulation2012). Other work links the PfC to reward-based decision-making during vicarious and personal video game play (MorelliVicariousPersonal2015) suggesting hidden potential for alignment with agent evaluation metrics. Therefore, we propose mapping fNIRS signals to various degrees of agent performance, a first step toward creating an fNIRS-driven RLHF system.
@@
We formalize the structure of the complete dataset as $\mathcal{D}$, and the subset of data from each participant $k$, denoted as $\mathcal{D}^{k}$. Dataset $\mathcal{D}^{k}$ further divides into two subsets, the neural data set and the task data set.
Neural Data Set: We define the neural dataset $\mathcal{N}^{k}$ as the set of neural channel recordings over time for some participant $k$. Let $\mathcal{N}_{i}\in\mathbb{R}^{M\times T}$ denote the neural signal data matrix at instance $i$, where $M$ is the number of neural channels and $T$ is the number of timestamps for some participant. A single neural data vector at timestamp $t$ across all channels can be denoted as $\mathbf{n}_{t}\in\mathbb{R}^{M}$. The signal at a specific timestamp $t$ and channel $m$ can be expressed as $n_{t,m}$, where $t\in\{1,\dots,T\}$ and $m\in\{1,\dots,M\}$.
Task Data Set: We define the Task Dataset $\mathcal{H}^{k}$ as the set of agent transition variables, or learning task statistics, over time for some participant $k$. Let $\mathcal{H}_{i}\in\mathbb{R}^{P\times T}$ represent the task data matrix at instance $i$, where $P$ is the number of task variables and $T$ is the number of timestamps. We may describe a vector of learning task statistics as $\mathcal{H}_{k}=\{S_{t},A_{t},R_{t},S_{t+1},V_{ct},B_{ct},E_{ct}\}_{t=0}^{t=T},$ where $S_{t}$ is the agent’s state at time $t$, $A_{t}$ is the action chosen by the agent or human, $R_{t}$ is the reward, $S_{t+1}$ is the next state, and $V_{t},B_{t},E_{t}$ represent various agent performance values. An instance of a task statistic data point at some timestamp for a specific task variable can be denoted as $h_{t,p}=f(H_{p},T_{t})$, where $p\in\{1,\dots,P\}$. After the raw neural data has been pre-processed, we propose that some relationship exists between features of the neural data and the learning task statistics outlined above, where $x$ is a neural feature vector. We let $\hat{y}=\phi(x_{i},h_{i})$ denote the relationship between a vector of human neural data features at some instance $i$ and its agent performance label.
@@
Participants: We recruited 25 participants for a mixed within and between-participants study. Participants were between 19-27 years old and were recruited through physical and virtual flyers. 14 participants were female and 11 were male. Each participant completed 3-4 conditions out of a possible six conditions, resulting in at least 10 participants per condition.
@@
High-Level Experimental Procedure: A complete experimental procedure included the administration an informed consent form, configuration the fNIRS device, instruction of tasks, and two post-experiment questionnaires (Figure 3).
Figure 4: Multi-Subject Classification Performance: This figure illustrates classification performance (F1) for binary and multi-class models. Binary models attempted to classify optimal vs. sub-optimal behavior, while multi-class models attempted to differentiate between optimal, sub-optimal or worst-case agent behavior.
Each task consisted of a set of episodes in which an agent attempts to complete a domain-specific goal. The participant interface with an agent was either passive or active. A condition is any combination of domain and interaction. Participants completed 3-4 conditions that lasted between 2-5 minutes each. Conditions were randomly selected, but passive tasks were completed first. Participants were seated 24 inches from the computer screen that displayed the agent and domain. A 20-second rest/pause was taken before each task to calibrate the fNIRS device. The participant completed a NASA-TLX questionnaire for each task, and a post-task questionnaire at the end of the study.
@@
$p^{\prime}=\begin{cases}0&\text{if }i=b\\ p_{w}+p_{b}&\text{if }i=w\\ p_{i}&\text{otherwise}\end{cases}$
@@
Optimal Lunar Lander agents often landed between the flags without crashing. Sub-optimal agents often landed outside of the goal location. Worst-case agents crashed. Optimal Flappy Bird agents often completed long episodes (15+ pipes cleared). Sub-optimal Flappy agents completed medium-length episodes (5-15 pipes). Worst-Case Flappy agents often completed short episodes ($\leq$ 5 pipes).
In an environment with a continuous action space (Robot Fetch and Place), sub-optimal and worst-case action selection was implemented differently. Sub-optimal actions altered the robot’s goal, resulting in fluid, swift movements to the wrong goal state. Worst-case actions did not change the robot’s goal, but added random variance to joint positions, resulting in unnatural, abrupt movements or dropping/throwing the block.
@@
Active Task: The active condition instructed a participant to physically guide an agent towards its goal using a keyboard or joystick. Participants used a basic computer keyboard to guide Flappy Bird and Lunar Lander agents, and an Xbox Controller to guide the Robot arm. As they played, the participant’s chosen actions, earned rewards and states of the environment were saved. The neural and task data were labeled and saved identically to Passive Tasks.
@@
Figure 6: Multi-Policy Agreement Heatmap: These three heatmaps visually illustrate the point of failure and the degree of failure within each optimality class. Cool colors indicate low error and warm colors are high.
Participant Data Summary: Two participant were excluded from data analysis due to technical difficulties or non-consent to the release of their neural data. Participants with inefficient data are also included in the dataset so researchers can explore the effect of noisy fNIRS data for future analyses (suboptData2024). A Participant Data Summary is included with the dataset outlining eligibility and noteworthy disturbances during data collection.
@@
Classification Labels: To train both classifiers and regressors, both discrete and continuous labels were recorded. These labels are equivalent to the calculated degree of failure and are assigned to all actions after the point of failure in an episode. Binary performance labels are denoted as $\mathbf{B}$. At some time step $t$, $\mathbf{B}_{t}$ may be some binary number from the set $B_{t}\in\{0,1\}$ where 0 represents optimal behavior and 11 represents sub-optimal behavior. Multi-class performance labels are denoted as $\mathbf{V}$. At some time step $t$, $\mathbf{V}_{t}$ may be some discrete number from the set $V_{t}\in\{0,1,2\}$ where 0 represents optimal, $1$ represents sub-optimal and $2$ represents worst-case behavior. Continuous performance errors can be denoted as $\mathbf{E}_{t}\in\mathbb{R}^{m}$. At some time step $t$, $\mathbf{E}_{t}$ may be some continuous value where lower values indicate near-optimal actions and higher values indicate sub-optimal to worst-case actions. This error value was designated by the multi-policy agreement system outlined below.
Multi-Policy Action Agreement: To calculate continuous performance labels, we designed a multi-policy action classification system. In any reinforcement learning problem, there may be more than one optimal path to the same goal. To avoid mislabeling agent behavior, we compare the agent’s chosen action with 10 near-optimal policies. An error value is calculated between the agent’s chosen action and each near-optimal policy’s action. Then the average is calculated across the $K$ policies. The error between the agent’s discrete action and the $k$-th near-optimal policy $\pi_{k}\left(s_{t}\right)$ is calculated using Kullback-Leibler (KL) Divergence:
@@
Single-Subject Cross-Participant Fine-Tuned Single-Subject Cross-Participant Fine-Tuned
@@
Table 1: Performance Across Multi-Subject Models (MLP): Average classifier performance (F1) for various levels of Model Granularity and Conditions. Multi-subject and binary cross-validation saw mostly insignificant performance, but fine-tuning these models increased binary and multi-class model classification by nearly 17% and 41%, respectively. Random chance F1 performance was 0.50 for binary and 0.33 for multi-class classification.
Figure 6 shows three heatmaps showing multi-policy action agreement over three episodes of various optimality classes. Each performance category is shown to be distinct to our system and a point of failure can be identified within an episode.
@@
Time Classification Approach: We use a sliding window approach that extracts a fixed duration of seconds, where each window was assigned a single label. In our experiments, window length varied per condition and were chosen through ablation studies. They were generally 5 to 7 seconds in length and 1 to 2 seconds in stride to address the $5$ to $7$ second latency of the fNIRS signal. The label assigned to an entire window was designated based on the endpoint label.
@@
Training Paradigms: Our models were trained and evaluated using one of three different paradigms: single-subject, multi-subject and fine-tuned training. These paradigms are based on prior work that aimed to better calibrate BCI models for specific users (huangfNIRS2MW2021). As illustrated in 3, a single-subject model is trained on data from one participant and is validated using withheld data from the same participant. Multi-subject models were trained on data from a set of participants and evaluated using withheld data from the same set. Fine-tuned models are multi-subject models that were further trained with a fraction of a target participant’s data and evaluated using withheld data from that target participant. Multi-subject and fine-tuned training data was exclusive to one condition, or a set of passive or active conditions.
@@
Figure 7: Regressors for Multi-Subject Robot Passive Task: This figure compares actual and predicted performance values for participants passively observing the Robot condition. The model uses fNIRS data to predict a continuous performance label. All models a showed strong correlation between fNIRS signals and agent performance, with MSE, $R^{2}$ values, and corresponding scatter plots.
Cross-Validation Studies: Each multi-subject model was examined using a leave-one-subject-out (LOPO) approach, sometimes two. From the set of participants in a condition set, one or two of those participants were completely withheld as a cross-validation testing group. These participants were excluded from the training set, allowing cross-participant evaluation of the model within the same condition.

### v3

# Towards Reinforcement Learning from Neural Feedback: Mapping fNIRS Signals to Agent Performance

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) is a methodology that aligns agent behavior with human preferences by integrating user feedback into the agent’s training process. This paper introduces a framework that guides agent training through implicit neural signals, with a focus on the neural classification problem. Our work presents and releases a novel dataset of functional near-infrared spectroscopy (fNIRS) recordings collected from 25 human participants across three domains: Pick-and-Place Robot, Lunar Lander, and Flappy Bird. We train multiple classifiers to predict varying levels of agent performance (optimal, suboptimal, or worst-case) from windows of preprocessed fNIRS features, achieving an average F1 score of 67% for binary and 46% for multi-class classification across conditions and domains. We also train multiple regressors to predict the degree of deviation between an agent’s chosen action and a set of near-optimal policy actions, providing a continuous measure of performance. Finally, we evaluate cross-subject generalization and show that fine-tuning pre-trained models with a small sample of subject-specific data increases average F1 scores by 17% and 41% for binary and multi-class models, respectively. Our results demonstrate that mapping implicit fNIRS signals to agent performance is feasible and can be improved, laying the foundation for future Reinforcement Learning from Neural Feedback (RLNF) systems.
@@
Codehttps://github.com/your-profile/NeuroLoop-Classification/tree/aaai26
@@ Introduction
Recent advances in Reinforcement Learning from Human Feedback (RLHF) have become crucial for training and fine-tuning state-of-the-art systems (Wu et al. 2023; Mosqueira-Rey et al. 2022). Specifically, RLHF has seen significant success by addressing limitations that plague common Reinforcement Learning (RL) techniques. Integrating human feedback has been incredibly beneficial for aligning agent behavior with human preferences. However, many feedback methods require active participation and/or expert demonstration for evaluative feedback to be most effective (Christiano et al. 2017; Li et al. 2019; Retzlaff et al. 2024). In addition, most approaches are limited to explicit feedback, which may lack insight into the nuances of human decision-making and internal assessments. Techniques that retrieve richer feedback may require tasks that impose greater cognitive effort on users (Lindner and El-Assady 2022). These drawbacks can be inconvenient for the evaluator or limit accessibility to certain user groups. They may also contribute to higher mental workload and unnatural, shallow feedback (Casper et al. 2023). These limitations may make it difficult to develop adaptive AI technologies that fully align with human preferences while being available to all user groups.
Figure 1: NEURO-LOOP: A high-level diagram of the proposed framework. This paper addresses the neural classification problem with the intent to apply trained models to the proposed pipeline.
Implicit human feedback systems present a valuable alternative by allowing agents to learn from natural human responses. Modern methods focus on facial expression or gesture classification to adapt agent behavior at little to no additional cognitive cost to the teacher (Cui et al. 2021). However, some of these procedures still require explicit instruction and conscious physical effort to adjust gestures and expressions. Further, reliance on private training data poses challenges for data collection and distribution.
To address these limitations, we propose NEURO-LOOP, a fully implicit neural feedback framework that leverages fNIRS signals to directly guide agents’ training using RLHF techniques. This framework employs passive Brain-Computer Interfaces (BCI) to align agent behavior with user intent through passive observation. We employ a functional near-infrared spectroscopy (fNIRS) device to record brain activity. fNIRS is a non-invasive neuro-imaging tool that can track sustained cognitive states such as mental workload, attention, and decision-making. This technology measures the hemodynamic response, or change in blood flow, in the prefrontal cortex (PFC) over time.
This paper addresses the neural classification problem as the first step toward integrating implicit fNIRS signals into RLHF frameworks. Our contributions are outlined below:
@@
  • We demonstrate that machine learning can distinguish multiple levels of agent performance, extending past binary classification and enabling finer-grained feedback while maintaining low cognitive workload for users.
  • We evaluate model performance and generalization between participants, suggesting that cross-subject transfer remains a challenge, but subject-specific calibration using limited data is feasible and improves performance.
We further compare model performance and user workload across both passive (observation) and active (physical demonstration) tasks, establishing a benchmark for future research in Reinforcement Learning from Neural Feedback (RLNF).
@@
Reinforcement Learning from Human Feedback (RLHF) allows agents to learn from human preferences through a range of approaches. These methods can be described as having either implicit or explicit feedback mechanisms.
@@
Explicit Feedback: Explicit feedback is the most common modality used to shape an agent’s policy to align with human expectations. Methods in this category often provide binary or scalar evaluative signals, such as preference labels or human-provided rewards, to shape the agent’s policy or learning process (Griffith et al. 2013; Knox and Stone 2011). Past research indicates that evaluative metrics with wider ranges have greater training success (Yu et al. 2023). These interactions are often shorter-term due to increased cognitive workload and sustained physical demand. Imitation learning is a related approach that often requires expert demonstration or active participation for the agent to learn from user examples effectively (Hussein et al. 2017).
@@
Implicit Feedback: Implicit feedback leverages unobtrusive human responses to enhance reinforcement learning agent behavior. These feedback techniques often include gaze tracking (Veeriah et al. 2016), expression recognition (Arakawa et al. 2018), or gesture classification (Cui et al. 2021) to shape agent policies. One drawback is that some of these procedures require explicit instruction, as participants must consciously and physically adjust their gestures or expressions to provide meaningful feedback. We propose using neural signals to communicate evaluative feedback directly with minimal instruction.
@@
Figure 2: Setup: Participants sat 24 inches in front of a computer screen. The fNIRS device is a headband that shines pulsating infrared light into the PFC to detect changes in blood flow.
@@
Figure 3: Training Paradigms: Diagram of the three training paradigms and Experimental Protocol. (a) Single-subject models are trained on a set of data from one participant and are evaluated using withheld data from the same participant. (b) Multi-subject models are trained on a set of participants and are evaluated using withheld data from the same set. (c) Fine-tuned models are multi-subject models calibrated with a fraction of a target participant’s data.
@@
ErrPs for Human-Robot Interaction: Error-related potentials (ErrPs) are neural signals that occur when the brain perceives an error. These signals, recorded via EEG, have been explored as a method for improving human-robot interaction. Previous research has applied ErrP signals to intervene in robotic decision-making by providing a binary indicator of error or dissatisfaction. Some work has even integrated ErrPs into RLHF algorithms, demonstrating improved training efficiency (Vukelić et al. 2023; Agarwal et al. 2020; Xu et al. 2021). EEG signals are often transient and susceptible to artifacts (Rihet et al. 2024). We explore the use of fNIRS as a complementary signal for brain-driven RLHF systems. Compared to other neuro-imaging devices, fNIRS produces signals that offer more tolerance to physical movement, greater portability, and higher spatial resolution (Pinti et al. 2018). Using the fNIRS system described by Giles (Blaney et al. 2020), we collected phase and intensity signals, making fNIRS suitable for long-horizon naturalistic settings outside the laboratory.
@@
fNIRS for Human-Centered Technologies: Recent advancements have made it possible for brain signals to be used in human-centered technologies (Ayaz and Dehais 2021). When applied to adaptive frameworks, fNIRS data often monitors cognitive workload, allowing systems to adjust to user strain. Past work has explored applications like unmanned aerial vehicles (UAV) (Afergan et al. 2014), and human-robot systems (Solovey et al. 2012; Roy et al. 2020).
@@
Motivation: fNIRS is quite under-explored in reinforcement learning and human-in-the-loop machine learning. fNIRS also offers beneficial characteristics and features that EEG generally lacks. Prior research has correlated fNIRS signals to reward-based gameplay events like video game performance (Li et al. 2018) and physical game performance (Bronner et al. 2012). Other work links the PFC to reward-based decision-making during vicarious and personal video game play (Morelli et al. 2015), suggesting hidden potential for alignment with agent evaluation metrics. Therefore, we propose mapping fNIRS signals to various degrees of agent performance, a first step toward creating an fNIRS-driven RLHF system.
@@
We formalize the structure of the complete dataset as $\mathcal{D}$, and the subset of data from each participant $k$, denoted as $\mathcal{D}^{k}$. Dataset $\mathcal{D}^{k}$ is further divided into two subsets, the neural dataset and the task dataset.
Neural Dataset: We define the neural dataset $\mathcal{N}^{k}$ as the set of neural channel recordings over time for some participant $k$. Let $\mathcal{N}_{i}\in\mathbb{R}^{M\times T}$ denote the neural signal matrix at instance $i$, where $M$ is the number of neural channels and $T$ is the number of timestamps for some participant. A single neural data vector at timestamp $t$ across all channels can be denoted as $\mathbf{n}_{t}\in\mathbb{R}^{M}$. The signal at a specific timestamp $t$ and channel $m$ can be expressed as $n_{t,m}$, where $t\in\{0,\dots,T\}$ and $m\in\{0,\dots,M\}$.
Task Dataset: We define the Task dataset $\mathcal{H}^{k}$ as the set of agent transition variables, or learning task statistics, over time for some participant $k$. Let $\mathcal{H}_{i}\in\mathbb{R}^{P\times T}$ represent the task data matrix at instance $i$, where $P$ is the number of task variables and $T$ is the number of timestamps. We may describe a vector of learning task statistics as $\mathbf{h}_{k}=\{S_{t},A_{t},R_{t},S_{t+1},V_{t},B_{t},E_{t}\}_{t=0}^{t=T},$ where $S_{t}$ is the agent’s state at time $t$, $A_{t}$ is the action chosen by the agent or human, $R_{t}$ is the reward, $S_{t+1}$ is the next state, and $V_{t},B_{t},E_{t}$ represent various agent performance values. An instance of a task statistic data point at some timestamp for a specific task variable can be denoted as $h_{t,p}=f(H_{p},T_{t})$, where $p\in\{0,\dots,P\}$. After the raw neural data has been pre-processed, we hypothesize that a relationship exists between features of the neural data and the learning task statistics outlined above, where $x$ is a neural feature vector. We let $\hat{y}=\phi(x_{i},h_{i})$ denote the relationship between a vector of human neural data features at some instance $i$ and its agent performance label.
@@
Participants: We recruited 25 participants for a mixed within and between-participants study. Participants were between 19 to 27 years old and were recruited through physical and virtual flyers. Fourteen participants identified as female and eleven as male. Each participant completed 3-4 conditions out of a possible six conditions, resulting in at least 10 participants per condition.
@@
High-Level Experimental Procedure: A complete experimental procedure included the administration of an informed consent form, configuration of the fNIRS device, task instruction, and two post-experiment questionnaires (Figure 3).
Figure 4: Multi-Subject Classification Performance: This figure illustrates classification performance (F1) for binary and multi-class models. Binary models attempted to classify optimal vs. suboptimal behavior, while multi-class models attempted to differentiate between optimal, suboptimal or worst-case agent behavior.
Each task consisted of a set of episodes in which an agent attempts to complete a domain-specific goal. The participant’s interaction with the agent was either passive or active. A condition is any combination of domain and interaction. Participants completed 3-4 conditions that lasted between 2-5 minutes each. Conditions were randomly selected, but passive tasks were completed first. Participants were seated 24 inches from the computer screen that displayed the agent and domain. A 20-second rest/pause was taken before each task to calibrate the fNIRS device. The participant completed a NASA-TLX questionnaire for each task, and a post-task questionnaire at the end of the study.
@@
$p^{\prime}_{i}=\begin{cases}0&\text{if }i=b\\ p_{w}+p_{b}&\text{if }i=w\\ p_{i}&\text{otherwise}\end{cases}$
@@
Optimal Lunar Lander agents often landed between the flags without crashing. Suboptimal agents often landed outside of the goal location. Worst-case agents crashed. Optimal Flappy Bird agents often completed long episodes (15+ pipes cleared). Suboptimal Flappy agents completed medium-length episodes (5-15 pipes). Worst-case Flappy agents often completed short episodes ($\leq$ 5 pipes).
In an environment with a continuous action space (Robot Fetch and Place), suboptimal and worst-case action selection was implemented differently. Suboptimal actions altered the robot’s goal, resulting in fluid, swift movements to the wrong goal state. Worst-case actions did not change the robot’s goal, but added random variance to joint positions, resulting in unnatural, abrupt movements or dropping/throwing the block.
@@
Active Task: The active condition instructed a participant to physically guide an agent towards its goal using a keyboard or joystick. Participants used a basic computer keyboard to guide Flappy Bird and Lunar Lander agents, and an Xbox Controller to guide the Robot arm. As they played, the participant’s chosen actions, earned rewards, and seen states were saved. The neural and task data were labeled and saved identically to Passive Tasks.
@@
Figure 6: Multi-Policy Agreement Heatmap: These three heatmaps visually illustrate the point of failure and the degree of failure within each optimality class. Cool colors indicate low error, while warm colors indicate high error.
Participant Data Summary: Two participants were excluded from data analysis due to technical difficulties or non-consent to the release of their neural data. Participants with inefficient data are also included in the dataset so researchers can explore the effect of noisy fNIRS data for future analyses (Muslimani and Taylor 2024). A Participant Data Summary is included with the dataset outlining eligibility and noteworthy disturbances during data collection.
@@
Classification Labels: To train classifiers and regressors, both discrete and continuous labels were recorded. These labels are equivalent to the calculated degree of failure and are assigned to all actions after the point of failure in an episode. Binary performance labels are denoted as $B_{t}\in\{0,1\}$, where 0 represents optimal behavior and $1$ represents suboptimal behavior. Multi-class performance labels are denoted as $\mathbf{V}$. At some time step $t$, $\mathbf{V}_{t}$ may be some discrete number from the set $V_{t}\in\{0,1,2\}$ where 0 represents optimal, $1$ represents suboptimal, and $2$ represents worst-case behavior. Continuous performance errors are denoted as $E(t)\in\mathbb{R}$. At some time step $t$, $\mathbf{E}_{t}$ may be some continuous value where lower values indicate near-optimal actions and higher values indicate suboptimal to worst-case actions. This error value was designated by the multi-policy agreement system outlined below.
Multi-Policy Action Agreement: To calculate continuous performance labels, we designed a multi-policy action classification system. In any reinforcement learning problem, there may be more than one optimal path to the same goal. To avoid mislabeling agent behavior, we compare the agent’s chosen action with $K$ near-optimal policies, where $K=10$. An error value is calculated between the agent’s chosen action and each near-optimal policy’s action. Then the average is calculated across the $K$ policies. The error between the agent’s discrete action and the $k$-th near-optimal policy $\pi_{k}\left(s_{t}\right)$ is calculated using Kullback-Leibler (KL) Divergence:
@@
Single-Subject Cross-Subject Fine-Tuned Single-Subject Cross-Subject Fine-Tuned
@@
Table 1: Performance Across Multi-Subject Models (MLP): Average classifier performance (F1) for various levels of Model Granularity and Conditions. Binary and multi-class cross-validation saw mostly insignificant performance, but fine-tuning these models increased binary and multi-class model classification by nearly 17% and 41%, respectively. Random chance F1 performance was 0.50 for binary and 0.33 for multi-class classification.
Figure 6 shows three heatmaps illustrating multi-policy action agreement over three episodes for each optimality class. Each performance category is shown to be distinct to our system and a point of failure can be identified within an episode.
@@
Time Classification Approach: We use a sliding window approach that extracts fixed-duration time windows, where each window was assigned a single label. In our experiments, window length varied per condition and was chosen through parameter studies. They were generally 5 to 7 seconds in length and 1 to 2 seconds in stride to address the $5$ to $7$ second latency of the fNIRS signal. The label assigned to an entire window was designated based on the endpoint label.
@@
Training Paradigms: Our models were trained and evaluated using one of three different paradigms: single-subject, multi-subject and fine-tuned training. These paradigms are based on prior work that aimed to better calibrate BCI models for specific users (Huang et al. 2021). As illustrated in 3, a single-subject model is trained on data from one participant and is validated using withheld data from the same participant. Multi-subject models were trained on data from a set of participants and evaluated using withheld data from the same set. Fine-tuned models are multi-subject models that were further trained with a fraction of a target participant’s data and evaluated using withheld data from that target participant. Multi-subject and fine-tuned training data was exclusive to one condition, or a set of passive or active conditions.
@@
Figure 7: Regression for Multi-Subject Robot Passive Task: This figure compares actual and predicted performance values for participants passively observing the Robot condition. The model uses fNIRS data to predict a continuous performance label. All models showed strong correlation between fNIRS signals and agent performance, with MSE, $R^{2}$ values, and corresponding scatter plots.
Cross-Validation Studies: leave-one-subject-out (LOPO), and in some cases leave-two-subjects-out, approach. From the set of participants in a condition set, one or two of those participants were completely withheld as a cross-validation testing group. These participants were excluded from the training set, allowing cross-subject evaluation of the model within the same condition.